OpenStudy (anonymous):

Find the radius of the circle with equation x^2 + y^2 +8x + 8y + 28 =0 Please help I am so stuck

OpenStudy (anonymous):

So what you need to do is complete the square. Do this one at a time. Take the x2 and 8x and add (1/2)(8)^2=16 so you get (x+4)^2 do the same thing with y and get (y+4)^2 Then because you added 16 twice to complete the square twice, you need to subtract 16 twice. This -32 combines with the 28 to make -4 add -4 to other side and get 4 so there you have it the radius is 2 how do i know this? Because x^2+y^2=r^2 is the general form for a circle. Now that it is in the form you can pick out r which is on the other side of the equation as 2^2 therefore the radius is 2

OpenStudy (anonymous):

\[x^2+y^2+8x+8y+28=0\] \[x^2+y^2+8x+8y+16+16=-28+16+16\] \[(x^2+8x+16)+(y^2+8y+16)=4\] \[(x+4)^2+(y+4)^2=4\] Now we know that the eq. of circle with centre(h,k) is given by : \[(x-h)^2+(y-k)^2=r^2\] Comparing, h=-4 k=-4 r=2

OpenStudy (anonymous):

i have just completed the square .

OpenStudy (anonymous):

Thank you guys so much. It helped a lot.:)

OpenStudy (raden):

alternative : if given an eqution of circle : x^2+y^2+Ax+By+C=0, so its radius = r = sqrt (A^2/4 + B^2/4 - C ) r = sqrt (8^2/4 + 8^2/4 - 28 ) r = sqrt (16+16-28) r = sqrt(4) = 2